The length of a rectangular pool is 8 more feet then the width. The perimeter of the pool is 96 feet. What system of equations can be used to find the pools width and length?
Thank you:)

Answer by Cromlix(4381)

Perimeter = 2x length + 2 x width
=2(x + 8) + 2x
= 2x + 16 + 2x
= 4x + 16 = 96
= 4x = 80
= x = 20
Therefore length = 28 feet
width = 20 feet
Perimeter = 2(28) + 2(20)
= 56 + 40
= 96 feet.